OpenStudy (anonymous):

Points A and B lie on a circle centered at point O. If length of AB/radius= π/10, what is the ratio of the area of sector AOB to the area of the circle? A. 1/10 B. π/10 C. 1/20 D. π/20

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@BSwan

OpenStudy (anonymous):



OpenStudy (anonymous):

AB/radius= π/10, radius =\(\large \frac{10 AB}{\pi}\)

OpenStudy (anonymous):

area of circle = (radius\()^2 \pi\) = \(\large( \frac{10 AB}{\pi})^2 \pi\) = \(\large \frac{(10 AB)^2 }{\pi} \)

OpenStudy (anonymous):

nw how to find area of sector :D

OpenStudy (anonymous):

got it. its 1/20. Thanks ur the best like always xD

OpenStudy (anonymous):

aww ok :)

OpenStudy (anonymous):

YW :)

OpenStudy (anonymous):

:DDDDD